# Restyling Unsupervised Concept Based Interpretable Networks with Generative Models

### v1

@@ 3.4 Training losses
Based on the previous discussion about concept based networks and our reconstruction pipeline design inspired from GAN inversion architectures, we define here our training loss $\mathcal{L}_{train}$ and each of its constitutive terms.
  • To verify the fidelity to output property, we define an output fidelity loss $\mathcal{L}_{of}$, which grants predictive capabilites to $g$. It is defined as the generalized cross entropy (CE) loss between $g$ and $f$:
@@
  • The most critical part of our training loss is the reconstruction loss $\mathcal{L}_{rec}^{G}$ computed through the pretrained generative model $G$, that gathers all constraints between inputs $x$ and their reconstruction $\tilde{x}=G(\Omega(\Phi(x),\Phi^{\prime}(x)))$. It combines $\ell_{1}$ and $\ell_{2}$ penalties, enforcing pixel-wise reconstruction for fidelity to input, with perceptual similarity LPIPS [82] and a final reconstruction classification term, both linked to viewability. The reconstruction classification term, defined as $CE(f(\tilde{x}),f(x))$, encourages the generative model to reconstruct $\tilde{x}$ with more classification specific features pertaining to input $x$. Similar losses have been introduced for inversion with StyleGAN and its training for post-hoc interpretation [43]. Our reconstruction loss is thus defined as follows:
@@
  • We impose the sparsity property along with two other regularizations, combined under the term $\mathcal{L}_{reg}$. More specifically, $\Psi$ is regularized to encourage sparsity of activations in $\Phi(x)$ through an $\ell_{1}$ penalty, and diversity while reducing redundancy in learnt dictionary $\Phi$ with a kernel orthogonality loss $\mathcal{L}_{orth}$, applied on weights of final convolution layer of $\Psi$ [74, 71]. Then, $\Omega$ is encouraged to predict latent vectors close to average latent vector $\bar{w}$, a common practice in GAN-inversion systems [68]. The regularization terms are written as follows:
@@
    $\mathcal{L}_{reg-\Omega}(x;\Omega)=||w_{x}^{+}-\bar{w}||_{2}^{2},\quad\mathcal{L}_{reg-\Psi}(x;\Psi)=\delta||\Phi(x)||_{1}+\mathcal{L}_{orth}(\Psi).$
@@
In the above equations, the loss hyperparameters are denoted by $\alpha,\beta,\gamma,\delta$. During training, the overall training loss is simultaneously optimized w.r.t parameters of only $\Psi,\Theta$ and $\Omega$, keeping $f$ and $G$ fixed.
@@ 3.5 Interpretation phase
We now describe the interpretation generation process, which can be divided in two parts. (1) Concept relevance estimation, that requires estimating the importance of any given concept function $\phi_{k}$ in prediction for a particular sample $x$ (local interpretation) or a class $c$ in general (global interpretation), and (2) Concept visualization, which pertains to visualizing the concept encoded by any given concept function $\phi_{k}$. We describe each of them in greater detail below:
(1) Concept relevance: Since our $g(x)$ adheres to structure of CoINs and among them closest to [51], the first step of relevance estimation almost follows as is. The estimation is based on concept activations $\Phi(x)$, and how the pooled version of $\Phi(x)$ is combined by the fully connected layer in $\Theta$ (with weights $\Theta_{W}$) to obtain the output logits. Note that this step does not rely on using the decoder/generator $G$. Specifically, the local relevance $r_{k}(x)$ of a concept function $\phi_{k}$ for a given sample $x$ is computed as the normalized version (between $[-1,1]$) of its contribution to logit of the predicted class $\hat{c}=g(x)$. The global relevance of concept function $\phi_{k}$ for a given class $c$, denoted $r_{k,c}$, is computed as the average of local relevance $r_{k}(x)$ for samples from class $c$. The above description is summarized in equation below wherein $\Theta_{W}^{k,\hat{c}}$ denotes the weight on concept $k$ for predicted class $\hat{c}$ in weight matrix $\Theta_{W}$:
@@
(2) Concept visualization: Once the importance of a concept function is estimated, one can extract the most important concepts for a sample $x$ or class $c$ by thresholding $r_{k}(x)$ or $r_{k,c}$ respectively. For visualizing any concept $\phi_{k}$, following previous CoINs, one can start by selecting most activating training samples for $\phi_{k}(x)$ over the whole training set or separately for each class it is highly relevant for. However, the core of our concept visualization process, is to utilize generator $G$ to visualize the impact of $\phi_{k}$ on any input $x$ it is relevant for. We do so by (1) directly modifying activation of $\phi_{k}(x)$ by a factor $\lambda\times\phi_{k}(x),\lambda\geq 0$ while keeping all other activations in $\Phi(x)$ intact, and (2) Visualizing generated output for increasing value of $\lambda$. The case of $\lambda=1$ corresponds to $\tilde{x}$, the reconstructed version of $x$. This process is summarized in Fig. 3. Extremely high values of $\lambda$ can push the predicted latent vectors far from the average latent vector in which case the generated output is less reliable. For our experiments, qualitatively we found maximum $\lambda$ upto 3 or 4 reliable with consistent modifications.
Figure 3: Overview of visualization for a given image $x$ and concept function $\phi_{k}$. By imputing a higher activation for $\phi_{k}(x)$ in $\Phi(x)$ (by a factor $\lambda=4$ in the figure), and comparing the obtained visualization to the original reconstruction $\tilde{x}$ (obtained with the untouched $\Phi(x)$), we interpret information encoded by $\phi_{k}$ about image $x$. For simplicity, we omit $\Phi^{\prime}(x)$ in the figure, as it remains the same.
@@ 4 Experiments and Results
Datasets . We experiment on image recognition tasks for large-scale images in three different domains with a greater focus on multi-class classification tasks: (1) Binary age classification (young/old) on CelebA-HQ [30], (2) fine-grained bird classification for 200 classes on Caltech-UCSD-Birds-200 (CUB-200) [70], and (3) fine-grained car model classification of 196 classes on Stanford Cars[41].
Implementation details . We use a ResNet50 [26] as our architecture for $f$. All images are processed at resolution $256\times 256$. For the pretrained GAN, on CelebA-HQ we use checkpoint released by NVIDIA [31]. On CUB-200 and Stanford Cars, we fine-tune an ImageNet [24] and LSUN Cars [79] checkpoint respectively to obtain the pretrained $G$ in a resource efficient manner. The fine-tuning is performed without any class-labels. We use a dictionary size $K=64$ on CelebA-HQ and $K=256$ on CUB-200 and Stanford-Cars. All experiments were conducted on a single V100-32GB GPU.We will release our code publicly upon publication. Complete details about architecture, training and evaluation are in appendix 0.B.
@@ 4.1 Evaluation strategy
One major goal of by-design interpretable architectures is to obtain high prediction performance. Thus, prediction accuracy of $g$ is the first metric we evaluate. We next discuss multiple functionally-grounded metrics [17] that evaluate the learnt concept dictionary $\Phi$ from an interpretability perspective and its use in visualization, including two novel metrics in the context of evaluating CoINs (“faithfulness” and “consistency”).
@@
Fidelity of reconstruction . Since reconstructed output plays a crucial role in our visualization pipeline, we evaluate how well does $G$ reconstruct the input. We do so by computing averaged per-sample mean squared error (MSE), perceptual distance (LPIPS) [82] and distance of overall distributions (FID) [28] of reconstructed images $\tilde{x}$ and original input images $x$.
@@
Faithfulness of concept dictionary $\Phi$ The aspect of faithfulness for a generic interpretation method asks the question “are the features identified in the interpretation truly relevant for the prediction process?” [7, 52]. This is generally computed via simulating “feature removal” from the input and observing the change in predictor’s output [27]. Simulating feature removal from input is relatively straightforward for saliency methods compared to concept-based methods, for example by setting the pixel value to 0. For CoIN systems, this is significantly more tricky, as concept activations $\phi_{k}(x)$ don’t represent the input $x$ exactly. However, through the decoder, we can evaluate if the concepts identified as relevant for an input encode information that is important for prediction. We adopt an approach similar to a previous proposal of faithfulness evaluation for audio interpretation systems [52]. Concretely, for a given sample $x$ with activation $\Phi(x)$, predicted class $\hat{c}$ and a threshold $\tau$, we first manipulate $\Phi(x)$ such that $\phi_{k}(x)$ is set to 0 if $r_{k}(x)>\tau,\forall k\in\{1,...,K\}$. That is, we “remove” all concepts with relevance greater than some threshold. This modified version of $\Phi(x)$ is referred to as $\Phi_{rem}(x)$. To compute faithfulness for a given $x$, denoted by $\text{FF}_{x}$, we compute the change in probability of the predicted class from original reconstructed sample $\tilde{x}=G(\Omega(\Phi(x),\Phi^{\prime}(x)))$ to new sample $x_{rem}=G(\Omega(\Phi_{rem}(x),\Phi^{\prime}(x)))$, that is, $\text{FF}_{x}=g(\tilde{x})_{\hat{c}}-g(x_{rem})_{\hat{c}}$. Ideally, we expect to see a drop in probability ($\text{FF}_{x}>0$) if the set of relevant concepts “truly” encode information relevant for classification. Following [52] we report the median of $\text{FF}_{x}$ over the test data for different thresholds $10<\tau<$ 1.
@@
Consistency of visualization . We expect during visualization of a given concept $\phi_{k}$ that a user observes similar semantic modifications across different images. Thus, we hypothesize that if modifying any specific concept activation $\phi_{k}(x)$ leads to consistent changes for different samples $x$, then generated output for two versions of $\Phi(x)$, one with $\phi_{k}(x)$ set to a large value and one with $\phi_{k}(x)=0$, should be separable in the embedding space of $f$ (all other concept activations unchanged). In other words, embeddings for images with high $\phi_{k}(x)$ and low $\phi_{k}(x)$ should be well separated. To compute this metric, we first create a dataset of generated images with two different sets of activations. For each of training and test set, this is done by first selecting a set of $N_{cc}$ samples for which $\phi_{k}$ is highly activating and relevant for. Then we find its maximum activation $\phi_{k}^{max}$ among these samples, and create two generated outputs for each of $N_{cc}$ samples, one $x_{k}^{+}$ such that $\phi_{k}(x_{k}^{+})$ is set to $\lambda\phi_{k}^{max}$ with $\lambda\geq 1$, and the other $x_{k}^{-}$ such that $\phi_{k}(x_{k}^{-})=0$. The two sets of generated images are then gathered into a single dataset $\mathcal{S}_{k}=\{(x_{k}^{+},1),(x_{k}^{-},0)\}$ such that $|\mathcal{S}_{k}|=2N_{cc}$, and we learn a binary classifier $\varphi_{k}:\mathcal{X}\rightarrow\{0,1\}$, from pooled feature maps of intermediate embedding of $f$. We train for binary classification for sets created from training data $\mathcal{S}_{k}^{\text{train}}$ and test on sets created from test data $\mathcal{S}_{k}^{\text{test}}$. Our concept consistency metric $CC_{k}$ for a given concept $k$ is thus obtained as the accuracy of the binary classifier on $\mathcal{S}_{k}^{\text{test}}$:
@@
Baselines . The primary comparison methods for us are CoINs that learn unsupervised concepts efficiently, even for large-scale images, FLINT [51] and FLAEM [59]. SENN [7] suffers from computational issues for large-scale images as it requires to compute jacobian of concept dictionary w.r.t input pixels for its loss computation. Thus, for all the metrics we compare with FLINT and FLAEM as our primary baselines. Additionally, for accuracy evaluation, we track the performance of our pretrained classifier $f$. Note that $f$ (ResNet50) is not an interpretable model and trained entirely for accuracy. Lastly, for faithfulness we compare with a “random” baseline that randomly selects concepts for whom activation is set to 0. Since there is no notion of threshold in selection, in order to make it comparable for a given threshold, we select the same number of concepts randomly as we would for our method. Our proposed system for all evaluations is abbreviated as VisCoIN (Visualizable CoIN).
@@ 4.2 Results and discussion
Quantitative results . Tab. 1(a) reports the test accuracy of all the evaluated systems. Our proposed system, VisCoIN performs competitively with the pretrained $f$ considered uninterpretable and purely trained for performance. It also performs better than the other recent CoIN systems for more complex classification tasks (CUB-200 and Stanford Cars) with large number of classes and diverse images. Metrics quantifying the fidelity of reconstruction on test data are in Tab. 2(a). The other baselines only optimize for pixel-wise reconstruction and FLINT achieves a lower MSE than VisCoIN. However, crucially, reconstruction from our method approximates the input data considerably better, in terms of perceptual similarity (LPIPS) and overall distribution (FID), which highly contributes to better viewability. Tab. 2(b) tabulates the median faithfulness $\text{FF}_{x}$ for the evaluated systems on 1000 random test samples for different thresholds. The performance of Random baseline being close to 0 even for small thresholds indicates that a random selection of concepts often contains little information relevant for classification of the predicted class. In contrast, concepts identified relevant as part of $g$ in VisCoIN tend to encode information about input that noticeably affects classification. In regard to other CoINs, while the faithfulness results are competitive on CelebA-HQ, for more complex datasets, concept dictionary in VisCoIN is significantly more faithful than FLINT or FLAEM, which do not demonstrate more faithfulness than the Random baseline. Finally, the mean and standard deviation for consistency of all concepts is reported in Tab. 1(b) with $\lambda=2$. Concepts learnt with VisCoIN demonstrate a higher mean consistency compared to baselines. The deviation across concepts is also lower for our method. We also evaluate consistency with higher values of $\lambda$ and observe increased separation with better classification performance (Appendix 0.C).
Table 1: (a) Accuracy (in %) of interpreter $g$ of CoIN systems, and of the baseline pretrained classifier $f$. (b) Mean and standard deviation for consistency $CC_{k}$ over all concept functions $\phi_{k}$ (binary accuracy in %).
Higher is better, the best performance is reported in
bold, second best in underline.
@@
Qualitative results . Fig. 4 shows visualization for different class-concept pairs across the three datasets that are determined to have high global relevance $r_{k,c}$ through predictive structure of $g(x)$, as described in Sec. 3.5. For each class-concept pair, we show two maximum activating training samples for the concept from the corresponding class, the reconstructed input from $\Phi(x)$ ($\lambda=1$) and the generated output with modified concept activation $\phi_{k}(x)$ by a factor $\lambda=4$. In all the illustrations, increasing the activation of the concept, i.e. moving from $\lambda=1$ to $4$, strongly emphasizes some specific concept in the generated output that can be clearly grounded to the input. For instance, increasing the activation of concept for “Red-eye” in Fig. 4(a) increases the size of red eye of the bird, a key feature of samples from class 25 (“Bronzed-cowbird”). We can also qualitatively verify that the reconstruction has a high enough quality that allows us to “view” the input sample through the generated output and ground the modifications in generated output to the input. However, we also observed that learnt concept functions can be prone to modifying more than one high-level feature in the image. For, e.g., in Fig. 4(d), increasing the concept activation increases both “eye-squint” and “beard” in the generated output. A longer discussion about this and other limitations is available in Appendix 0.F.
@@
Figure 4: Qualitative examples obtained for different concepts, classes on (a)-(b) CUB-200, (c)-(d) CelebA-HQ, (e)-(f) Stanford-Cars datasets. On each subfigure, first column corresponds to maximum activated samples $x$ for class-concept pairs with high relevance ($r_{k,c}>0.5$), second column to reconstructed image obtained with original $\Phi(x)$, and third column to the image obtained by imputing $4\times\phi_{k}(x)$ in $\Phi(x)$.
Ablation study . We ablate multiple aspects of our system, with detailed results in Appendix 0.D. Most notably, we observed a tradeoff induced by strength of reconstruction-classification loss (weight $\gamma$). A high $\gamma$ positively impacts the faithfulness but negatively impacts the perceptual similarity of reconstruction.

### v2

@@ 3.3 Training losses
Based on the previous discussion about concept-based networks and our proposed reconstruction pipeline design, we define here our training loss $\mathcal{L}_{train}$ and each of its constitutive terms.
  • For the fidelity to output property, we define an output fidelity loss $\mathcal{L}_{of}$, that grants predictive capabilities to $g$. It’s defined as generalized cross entropy (CE) between $g$ and $f$:
@@
  • The most critical part of our training loss is the reconstruction loss $\mathcal{L}_{rec}^{G}$ computed through the pretrained generative model $G$, that gathers all constraints between inputs $x$ and their reconstruction $\tilde{x}=G(\Omega(\Phi(x)))$. It combines $\ell_{1}$ and $\ell_{2}$ penalties, enforcing pixel-wise reconstruction for fidelity to input, with perceptual similarity LPIPS (Zhang et al., 2018b) and a final reconstruction classification term, both linked to viewability. The reconstruction classification term, defined as $CE(f(\tilde{x}),f(x))$, encourages the generative model to reconstruct $\tilde{x}$ with more classification specific features pertaining to input $x$. Similar losses have been introduced for inversion in generative model and its training for post-hoc interpretation (Lang et al., 2021). Our reconstruction loss is thus defined as follows:
@@
  • We impose the sparsity property along with two other regularizations, combined under the term $\mathcal{L}_{reg}$. More specifically, $\Psi$ is regularized to encourage sparsity of activations in $\Phi(x)$ through an $\ell_{1}$ penalty, and diversity while reducing redundancy in learnt dictionary $\Phi$ with a kernel orthogonality loss $\mathcal{L}_{orth}$, applied on weights of final convolution layer of $\Psi$ (Xie et al., 2017; Wang et al., 2020). Then, $\Omega$ is encouraged to predict latent vectors close to average latent vector $\bar{w}$, a common practice in inversion systems of generative models (Tov et al., 2021). The regularization terms are written as follows:
@@
    $\mathcal{L}_{reg-\Omega}(x;\Omega)=||w_{x}-\bar{w}||_{2}^{2},\quad\mathcal{L}_{reg-\Psi}(x;\Psi)=\delta||\Phi(x)||_{1}+\mathcal{L}_{orth}(\Psi).$
@@
In the above equations, the loss hyperparameters are denoted by $\alpha,\beta,\gamma,\delta$. During training, $\mathcal{L}_{train}$ is simultaneously optimized w.r.t parameters of $\Psi,\Theta$ and $\Omega$, while keeping $f$ and $G$ fixed.
@@ 3.4 Interpretation phase
The interpretation generation process can be divided in two parts. (1) Concept relevance estimation, that requires estimating the importance of any given concept function $\phi_{k}$ in prediction for a particular sample $x$ (local interpretation) or a class $c$ in general (global interpretation), and (2) Concept visualization, which pertains to visualizing the concept encoded by any given concept function $\phi_{k}$. We describe each of them in greater detail below:
(1) Concept relevance: Since our $g(x)$ adheres to structure of CoINs and among them closest to Parekh et al. (2021), the first step of relevance estimation almost follows as is. The estimation is based on concept activations $\Phi(x)$, and how the pooled version of $\Phi(x)$ is combined by the fully connected layer in $\Theta$ (with weights $\Theta_{W}$) to obtain the output logits. Note that this step does not rely on using the decoder/generator $G$. Specifically, the local relevance $r_{k}(x)$ of a concept function $\phi_{k}$ for a given sample $x$ is computed as the normalized version (between $[-1,1]$) of its contribution to logit of the predicted class $\hat{c}=g(x)$. The global relevance of concept function $\phi_{k}$ for a given class $c$, denoted $r_{k,c}$, is computed as the average of local relevance $r_{k}(x)$ for samples from class $c$. The above description is summarized in equation below wherein $\Theta_{W}^{k,\hat{c}}$ denotes the weight on concept $k$ for predicted class $\hat{c}$ in weight matrix $\Theta_{W}$:
@@
(2) Concept visualization: Once the importance of a concept function is estimated, one can extract the most important concepts for a sample $x$ or class $c$ by thresholding $r_{k}(x)$ or $r_{k,c}$ respectively. For visualizing any concept $\phi_{k}$, following previous CoINs, one can start by selecting most activating training samples for $\phi_{k}(x)$ over the whole training set or separately for each class it is highly relevant for. However, the core of our concept visualization process, is to utilize generator $G$ to visualize the impact of $\phi_{k}$ on any input $x$ it is relevant for. We do so by (1) directly modifying activation of $\phi_{k}(x)$ by a factor $\lambda\times\phi_{k}(x),\lambda\geq 0$ while keeping all other activations in $\Phi(x)$ intact, and (2) Visualizing generated output for increasing value of $\lambda$. The case of $\lambda=1$ corresponds to $\tilde{x}$, the reconstructed version of $x$. This process is summarized in Fig. 3. Extremely high $\lambda$ can push the predicted latent vectors far from the average latent vector in which case the generated output is less reliable. For our experiments, qualitatively we found $\lambda$ upto 3 or 4 reliable with consistent modifications.
Figure 3: Visualization for a given image $x$ and concept function $\phi_{k}$. By imputing a higher activation for $\phi_{k}(x)$ in $\Phi(x)$ (by a factor $\lambda=4$ in the figure), and comparing the obtained visualization to the original reconstruction $\tilde{x}$ (obtained with the untouched $\Phi(x)$), we interpret information encoded by $\phi_{k}$ about image $x$.
@@ 4 Experiments and Results
Datasets . We experiment on image recognition tasks for large-scale images in three different domains with a greater focus on multi-class classification tasks: (1) Binary age classification (young/old) on CelebA-HQ (Karras et al., 2018), (2) fine-grained bird classification for 200 classes on Caltech-UCSD-Birds-200 (CUB-200) (Wah et al., 2011), and (3) fine-grained car model classification of 196 classes on Stanford Cars (Krause et al., 2013).
Implementation details . Experiments in main text use a ResNet50 (He et al., 2016) as our architecture for $f$, and StyleGAN2-ADA (Karras et al., 2020) for $G$. Experiments with different architectures for $G$ (ProgressiveGAN (Karras et al., 2018), $\beta$-VAE (Higgins et al., 2017) on FashionMNIST (Xiao et al., 2017)) and $f$ (ResNet101) can be found in Appendix D. All images are processed at resolution $256\times 256$. We use a dictionary size $K=64$ on CelebA-HQ and $K=256$ on CUB-200 and Stanford-Cars. All experiments were conducted on a single V100-32GB GPU. Complete details about network architectures, obtaining pretrained checkpoints for $f,G$, VisCoIN training and evaluation metric implementations are provided in Appendix C.
@@ 4.1 Evaluation strategy
One major goal of by-design interpretable architectures is to obtain high prediction performance. Prediction accuracy of $g$ is thus the first metric we evaluate. We next discuss multiple functionally-grounded metrics (Doshi-Velez & Kim, 2017) that evaluate the learnt concept dictionary $\Phi$ from an interpretability perspective and its use in visualization, including two novel metrics in the context of evaluating CoINs (“faithfulness” and “consistency”).
@@
Fidelity of reconstruction . Since reconstructed output plays a crucial role in our visualization pipeline, we evaluate how well does $G$ reconstruct the input. We compute averaged per-sample mean squared error (MSE), perceptual distance (LPIPS) (Zhang et al., 2018b) and distance of overall distributions (FID) (Heusel et al., 2017) of reconstructed images $\tilde{x}$ and original input images $x$.
@@
Faithfulness of concept dictionary $\Phi$ The aspect of faithfulness for a generic interpretation method asks the question “are the features identified in the interpretation truly relevant for the prediction process?” (Alvarez-Melis & Jaakkola, 2018a; Parekh et al., 2022). This is generally computed via simulating “feature removal” from the input and observing the change in predictor’s output (Hedström et al., 2023). Simulating feature removal from input is relatively straightforward for saliency methods compared to concept-based methods, for example by setting the pixel value to 0. For CoIN systems, this is significantly more tricky, as concept activations $\phi_{k}(x)$ don’t represent the input $x$ exactly. However, through the decoder, we can evaluate if the concepts identified as relevant for an input encode information that is important for prediction. We adopt an approach similar to a previous proposal of faithfulness evaluation for audio interpretation systems (Parekh et al., 2022). Concretely, for a given sample $x$ with activation $\Phi(x)$, predicted class $\hat{c}$ and a threshold $\tau$, we first manipulate $\Phi(x)$ such that $\phi_{k}(x)$ is set to 0 if $r_{k}(x)>\tau,\forall k\in\{1,...,K\}$. That is, we “remove” all concepts with relevance greater than some threshold. This modified version of $\Phi(x)$ is referred to as $\Phi_{rem}(x)$. To compute faithfulness for a given $x$, denoted by $\text{FF}_{x}$, we compute the change in probability of the predicted class from original reconstructed sample $\tilde{x}=G(\Omega(\Phi(x)))$ to new sample $x_{rem}=G(\Omega(\Phi_{rem}(x))$, that is, $\text{FF}_{x}=g(\tilde{x})_{\hat{c}}-g(x_{rem})_{\hat{c}}$. Ideally, we expect to see a drop in probability ($\text{FF}_{x}>0$) if the set of relevant concepts “truly” encode information relevant for classification. Following Parekh et al. (2022) we report the median of $\text{FF}_{x}$ over the test data for different thresholds $10<\tau<$ 1.
@@
Consistency of concept visualization . We expect during visualization of a given concept $\phi_{k}$ that a user observes similar semantic modifications across different images. Thus, we hypothesize that if modifying any specific concept activation $\phi_{k}(x)$ leads to consistent changes for different samples $x$, then generated output for two versions of $\Phi(x)$, one with $\phi_{k}(x)$ set to a large value and one with $\phi_{k}(x)=0$, should be separable in the embedding space of $f$ (all other concept activations unchanged). In other words, embeddings for images with high $\phi_{k}(x)$ and low $\phi_{k}(x)$ should be well separated. To compute this metric, we first create a dataset of generated images with two different sets of activations. For each of training and test set, this is done by first selecting a set of $N_{cc}$ samples for which $\phi_{k}$ is highly activating and relevant for. Then we find its maximum activation $\phi_{k}^{max}$ among these samples, and create two generated outputs for each of $N_{cc}$ samples, one $x_{k}^{+}$ such that $\phi_{k}(x_{k}^{+})$ is set to $\lambda\phi_{k}^{max}$ with $\lambda\geq 1$, and the other $x_{k}^{-}$ such that $\phi_{k}(x_{k}^{-})=0$. The two sets of generated images are then gathered into a single dataset $\mathcal{S}_{k}=\{(x_{k}^{+},1),(x_{k}^{-},0)\}$ such that $|\mathcal{S}_{k}|=2N_{cc}$, and we learn a binary classifier $\varphi_{k}:\mathcal{X}\rightarrow\{0,1\}$, from pooled feature maps of intermediate embedding of $f$. We train for binary classification for sets created from training data $\mathcal{S}_{k}^{\text{train}}$ and test on sets created from test data $\mathcal{S}_{k}^{\text{test}}$. Our concept consistency metric $CC_{k}$ for a given concept $k$ is thus obtained as the accuracy of the binary classifier on $\mathcal{S}_{k}^{\text{test}}$:
@@
Baselines . The primary comparison methods for us are CoINs that learn unsupervised concepts efficiently, even for large-scale images, FLINT (Parekh et al., 2021) and FLAEM (Sarkar et al., 2022). SENN (Alvarez-Melis & Jaakkola, 2018a) suffers from computational issues for large-scale images as it requires to compute jacobian of concept dictionary w.r.t input pixels for its loss computation. Thus, for all the metrics we compare with FLINT and FLAEM as our primary baselines. Additionally, for accuracy evaluation, we track the performance of our pretrained classifier $f$. Note that $f$ (ResNet50) is not an interpretable model and trained entirely for accuracy. Lastly, for faithfulness we compare with a “random” baseline that randomly selects concepts for whom activation is set to 0. Since there is no notion of threshold in selection, in order to make it comparable for a given threshold, we select the same number of concepts randomly as we would for our method. Our proposed system for all evaluations is abbreviated as VisCoIN (Visualizable CoIN).
@@ 4.2 Results and discussion
Quantitative results
Table 1(a) reports the test accuracy of all the evaluated systems. Our proposed system, VisCoIN performs competitively with the pretrained $f$ considered uninterpretable and purely trained for performance. It also performs better than the other recent CoIN systems for more complex classification tasks (CUB-200 and Stanford Cars) with large number of classes and diverse images. Metrics quantifying the fidelity of reconstruction on test data are in Table 2(a). The other baselines only optimize for pixel-wise reconstruction and FLINT achieves a lower MSE than VisCoIN. However, crucially, reconstruction from our method approximates the input data considerably better, in terms of perceptual similarity (LPIPS) and overall distribution (FID), which highly contributes to better viewability. Table 2(b) tabulates the median faithfulness $\text{FF}_{x}$ for the evaluated systems on 1000 random test samples for different thresholds. The performance of Random baseline being close to 0 even for small thresholds indicates that a random selection of concepts often contains little information relevant for classification of the predicted class. In contrast, concepts identified relevant as part of $g$ in VisCoIN tend to encode information about input that noticeably affects classification. In regard to other CoINs, while the faithfulness results are competitive on CelebA-HQ, for more complex datasets, concept dictionary in VisCoIN is significantly more faithful than FLINT or FLAEM, which do not demonstrate more faithfulness than the Random baseline. Finally, the mean and standard deviation for visualization consistency of all concepts is reported in Table 1(b) with $\lambda=2$. Concepts learnt with VisCoIN demonstrate a higher mean consistency of visualization compared to baselines. The deviation across concepts is also lower for our method. We also evaluate concept consistency with higher values of $\lambda$ and observe increased separation with better classification performance (Appendix E).
Table 1: (a) Accuracy (in %) of interpreter $g$ of CoIN systems, and of the baseline pretrained classifier $f$. (b) Mean and standard deviation for consistency $CC_{k}$ over all concept functions $\phi_{k}$ (binary accuracy in %). Higher is better, the best performance is reported in bold, second best in underline.
@@
Qualitative results . Fig. 4 shows visualization for different class-concept pairs across the three datasets that are determined to have high global relevance $r_{k,c}$ through predictive structure of $g(x)$, as described in Section 3.4. For each class-concept pair, we show two maximum activating training samples for the concept from the corresponding class, the reconstructed input from $\Phi(x)$ ($\lambda=1$) and the generated output with modified concept activation $\phi_{k}(x)$ by a factor $\lambda=4$. In all the illustrations, increasing the activation of the concept, i.e. moving from $\lambda=1$ to $4$, strongly emphasizes some specific concept in the generated output that can be clearly grounded to the input, and the name of the concept is then manually inferred. For instance, increasing the activation of concept for “Red-eye” in Fig. 4(a) increases the size of red eye of the bird, a key feature of samples from class 25 (“Bronzed-cowbird”). We can also qualitatively verify that the reconstruction has a high enough quality that allows us to “view” the input sample through the generated output and ground the modifications in generated output to the input. However, we also observed that learnt concept functions can be prone to modifying more than one high-level feature in the image. For, e.g., in Fig. 4(d), increasing the concept activation increases both “eye-squint” and “beard” in the generated output. A longer discussion about limitations is available in Appendix H.
@@
Figure 4: Qualitative examples obtained for different concepts, classes on (a)-(b) CUB-200, (c)-(d) CelebA-HQ, (e)-(f) Stanford-Cars datasets. On each subfigure, first column corresponds to maximum activated samples $x$ for class-concept pairs with high relevance ($r_{k,c}>0.5$), second column to reconstructed image obtained with original $\Phi(x)$, and third column to the image obtained by imputing $4\times\phi_{k}(x)$ in $\Phi(x)$. Red boxes manually added to indicate key regions of modifications in generated images.
Ablation study . We ablate multiple aspects of our system, with detailed results in Appendix F. Notably, we observed a tradeoff induced by strength of reconstruction-classification loss (weight $\gamma$). A high $\gamma$ positively impacts faithfulness, but negatively impacts perceptual similarity of reconstruction.